6. The sphere with mass 200g moving at a speed of 2m / s, is hit by a second sphere with mass 800g, which is at rest. After the impact the velocity of the first sphere is directed from left to right and size 1 m / s

Find the magnitude and velocity of the second sphere after hitting the first sphere. Friction is not taken into account.


1
Expert's answer
2021-02-15T00:58:13-0500

From the law of conservation of linear momentum, we have:


"m_1v_{1i}+m_2v_{2i}=m_1v_{1f}+m_2v_{2f},""v_{2f}=\dfrac{m_1v_{1i}+m_2v_{2i}-m_1v_{1f}}{m_2},""v_{2f}=\dfrac{0.2\ kg\cdot2\ \dfrac{m}{s}+0-0.2\ kg\cdot1\ \dfrac{m}{s}}{0.8\ kg}=0.25\ \dfrac{m}{s}."

The sign plus means that the second sphere moves to the right.